Two frames are needed with the same perimeter. One frame is in the shape of a square and the other in the shape of a regular pentagon. Each side of the square is 7 inches longer than each side of the pentagon. Find the dimensions of each frame.
Answer by Theo(13342)

side of the square is 35 inches long.
side of the pentagon is 28 inches long.
-----
4 * 35 = 140
5 * 28 = 140
-------
here's how it was done.
-----
a = side of the square
b = side of the pentagon
4a = p
5b = p
both equal to p (perimeter) so they are equal to each other.
4a = 5b
side of the square is 7 inches longer than side of the pentagon.
a = b + 7
replace a with b + 7 in the equation to get:
4 * (b+7) = 5b
this becomes:
4b + 28 = 5b
this becomes:
b = 28
-----
if b = 28, then a = 28 + 7 = 35
a = 35
b = 28
-----
side of square = 35 inches
side of pentagon = 28 inches
-----